September 13th, 2019

Email - All servers are now functioning correctly.

Email - Email Services are degraded. Engineers are working on the issues.

June 17th, 2019

Email - All email servers are online. DNS has propagated and services are back to normal.

June 12th, 2019

Email - Mail servers are being migrated. Services may be intermittent while DNS catches up. We apologise for any inconvenience.

Broadband - Service is Online.

September 26th, 2018

Broadband - We can see that service was restored on the majority of circuits at approximately 1pm. We are still awaiting details from BT regarding the cause and fix for the issue. Therefore services should be considered "at risk".

Broadband - Customers connected via Columbo Node may be experiencing connectivity issues. Engineers are investigating. First results show it to be a BT issue with multiple ISP's being affected.

April 3rd, 2017

Broadband - No further issues over the weekend so we now consider this to be no longer at risk.

March 31st, 2017

Broadband - kingston.core - Connectivity has been restored however the site should still be considered at-risk until we receive official confirmation that the power outage has been resolved.

Broadband - kingston.core - We've received confirmation from the power provider in the region that there is a power problem affecting the area. Power is expected to be restored between 14:30 and 15:30. Further updates will follow.

Broadband - We are currently investigating a loss of connectivity to kingston.core. Engineers are already investigating. Further information will be provided as and when it becomes available.

December 1st, 2016

Broadband - Services are now fully restored.

Broadband - All services via Columbo Node appear to be restored but should be considered as at risk until Engineers have completed their investigation.

Broadband - Customers connected via Columbo Node may be experiencing connectivity issues. Engineers are investigating.

November 29th, 2016

Hosting - All Services are now back to normal

Hosting - Some websites may be unresponsive due to a DDOS attack. Engineers are working on it.

November 21st, 2016

Email - Message Queues have now returned to normal.

Email - A SPAM attack is currently taking place that is causing delays to Email Protection users emails.

November 17th, 2016

Email - Maintenance completed.

Email - Service is Online.

November 16th, 2016

Email - Maintenance is being carried out on POP3/IMAP/Webmail services. We do not expect any downtime but these services should be considered as at risk.

July 5th, 2016

Broadband - Issues appear resolved. We will continue to monitor the situtaion.

Broadband - The network is experiencing issues accessing certain services, causing delays or inability to complete requests. Engineers are investigating the issue, expect disruption for the immediate future. Updates to follow.

June 21st, 2016

Email - Connectivity issues are resolved, however service is considered at risk until furher notice.

Hosting - Connectivity issues are resolved, however service is considered at risk until furher notice.

Virtual Servers - Connectivity issues are resolved, however service is considered at risk until furher notice.

Email - DDOS attacks via the TOR network are causing packet loss and connectivity issues. Engineers are currently dealing with the issue.

Hosting - DDOS attacks via the TOR network are causing packet loss and connectivity issues. Engineers are currently dealing with the issue.

Virtual Servers - DDOS attacks via the TOR network are causing packet loss and connectivity issues. Engineers are currently dealing with the issue.

June 17th, 2016

Broadband - Service is Online.

June 16th, 2016

Broadband - Large patches from Microsoft and users streaming the football is causing congestion across the UK.

January 28th, 2016

Email - A server issues was found and corrected. Some emails may have been delayed. Email is now functioning correctly.

Email - We are currently investigating an issue that is affecting Email. Updates to follow.

December 18th, 2015

Email - Outbound Server Authentication has remained stable and the system is no longer considered at risk.

Email - Customers using outbound server authentication will have experienced issues this morning. The issues have now been fixed. Sorry for any inconvenience.

November 30th, 2015

Email - Email is now routing via alternate servers and email should flow as normal. We have identified the source of the blacklist and are working to prevent future occurrences.

Email - We have received reports that one of our servers is currently on a blacklist. This is most likely due to a compromised customer. We are in the process of routing mail via alternate paths.

November 12th, 2015

Email - All servers are now responding correctly.

Email - Outgoing Mail Server has been slow to respond and may have been causing timeouts when sending email. This issue has now been resolved.

April 20th, 2015

Email - We have rebooted the problematic server and it is now functioning as expectedly. We are looking in to the cause. No email has been lost and email clients should reattempt automatically or via "Send & Receive". We apologise for any inconvenience caused.

Email - One of our sending servers are experiencing problems. We are looking into the cause.

February 20th, 2015

Virtual Servers - The crashed web server has now been repaired and is running. No data was lost. We apologise for an inconvenience caused.

February 19th, 2015

Virtual Servers - One of our web servers has crashed, affecting some customer websites. We are working on a fix.

November 12th, 2014

Broadband - Service is now running normally. No issues reported.

Broadband - Normal service is reported as being restored, however is still being closely monitored. Any customers still experiencing issues are advised to power cycle their equipment. Anyone still experiencing issues with their internet connection after a power cycle should contact technical support.

Broadband - Engineer work is still ongoing, further disruption is expected as equipment is worked on. We will continue to post updates as and when we get them.

Broadband - Customers may experience their Internet connection drop during the engineer work, most likely those in the London area. Recommend course of action is to power down the router for 15 minutes and then reconnect.

Broadband - Engineers have arrived on site. We are expecting further updates shortly.

Broadband - Customers have also reported experiencing packet loss which is also under investigation. We are expecting updates from the engineers shortly.

Broadband - Customers who are connected will likely be experiencing reduced speeds until the issue is resolved. An engineer has been dispatched and we expected to have an update within the next few hours.

Broadband - There are issues with a Core router affecting multiple Internet Connections. Please reboot your equipment and stand by for further updates.

October 1st, 2014

Broadband - After investigation this issue was resolved after maintenance on the core network was completed.

September 30th, 2014

Broadband - We are aware of issues with internet connections, it is currently being addressed by engineers. If you experience issues please reboot your equipment.

September 15th, 2014

Email - The spam has been removed successfully and normal service should be resumed shortly. Further delays are not likely but still possible, for at least the next 24 hours.

Email - We are experiencing delays with some outgoing emails due to an influx of spam emails. We are working to restore normal delivery but there may a delay to certain mail providers.

September 8th, 2014

Email - Email levels are now back to normal. Apologies for any inconvenience caused.

Email - We detected an issue without our outgoing mail service at 9am this morning which required the server to be rebooted. This has created a minor backlog which will result in delays with outgoing messages and potentially bounced messages. If you receive a non delivery notification please resend your email.

June 2nd, 2014

Email - Normal service appears to be resumed. There are no reported delays on emails to Google. We will continue to monitor the situation.

May 30th, 2014

Email - It appears Google are having issues with their mail servers, resulting in some email sent to Google addressed recipients being delayed. We have reported the issue to Google and are awaiting a response.

April 15th, 2014

Broadband - Traffic levels appear to be back to normal levels. No further issues are expected.

April 14th, 2014

Broadband - Some customers may be experiencing issues with slow speeds. We are investigating this with our upstream provider and will update you when we receive more information.

April 8th, 2014

Email - Services have remained stable and this incident is considered closed.

Email - Services have been restarted. All services have been restored but should be considered as at risk while we investigate the cause.

Email - There is currently an issue with the email system. Engineers are investigating.

March 27th, 2014

Email - This maintenance has now complete. There should be no further delays.

Email - As per our earlier announcement email protection users may be experiencing delayed mail for the next hour during emergency maintenance.

March 4th, 2014

Broadband - The problematic DNS server now appears to be fixed and DNS queries are back to normal.

Broadband - We are aware of DNS issues affecting a small proportion of our Broadband customer's. There appears to be an issue with one of our upstream provider's DNS servers resulting in some access to websites timing out. Our upstream provider is aware of this and is working towards a fix.

February 23rd, 2014

Email - The scheduled maintenance on Email Protection is now complete. All email is now flowing through our servers.

Email - Scheduled maintenance for Email Protection is now underway. All customers whose email is filtered will not recieve mail until the maintenance is completed.

February 20th, 2014

Broadband - Our upstream provider has reported that this issue has now been resolved. If customers are still experiencing issues, please restart your equipment/router.

Broadband - Our backbone provider is currently experiencing issues preventing some customers from connecting to our service. This is most likely to happen if you either reboot or lose connection on your router. This appears to be a problem with BT equipment and it is being investigated as a matter of urgency. Further updates will follow as they become available.

January 22nd, 2014

Email - Email is now flowing through our system as expected.

January 21st, 2014

Email - Email queues are returning to normal levels, there may be slight delays in outgoing email whilst normal queue levels resume.

Email - Users may be experiencing delays in outgoing emails. This is due to a customer being compromised and sending out spam. We will update shortly.

January 17th, 2014

Broadband - Issue appears to be resolved. We are awaiting engineer reports.

Broadband - Some customers may be experiencing issues with slow speeds. Our engineers are investigating and we will update this page as soon as information comes available.

December 15th, 2013

Virtual Servers - Power is now restored. All Services are now restored. An investigation into why backup power failed will take place and changes will be made to ensure it does not happen again.

Virtual Servers - Power outage affected all Lawyers Online Services. Normal service will be resumed shortly.

November 1st, 2013

Email - All servers and queues have been operating without incident for over 24hrs. This incident is considered closed.

October 31st, 2013

Email - Please note if you are receiving bounces with the words "security" and "permissions" this will be down to your email client / server caching DNS. Please reboot your equipment to refresh the cache. If you still receive this message, please contact technical on 0845 388 67 94

Email - All email is passing through our system as expected without delays.

Email - Email is now passing through our system. Users may experience small delays in the next hour, after this the email system should resume normal operations. If you receive any bounce messages mentioning "reject" you will need to resend these. We appologise for any inconvenience caused and we will monitor this situation closely.

Email - DNS issues are causing delays and bounces. We are looking into this and hope to fix soon.

Email - Email queues were cleared and have been monitored closely. No delays or issues have been noted or reported. This incident is considered closed.

October 29th, 2013

Email - Customers may experience delays with outgoing mail due a customer server being compromised. They produced thousands of spam emails. We are clearing the system and aim to have normal service resumed as soon as possible.

October 21st, 2013

Email - All email queues and services back to normal.

Email - We have idenitified a hardware issue and replaced the equipment. System is online and processing the backlog and no mail will have been lost. We will provide another status update first thing in the morning.

October 20th, 2013

Email - We are experiencing issues with our mail filtering service causing delays in emails. We are investigating and will update shortly.

October 11th, 2013

Email - Service is Online.

October 9th, 2013

Email - UK Lawyers be aware of hacked legal websites. There has been a run of emails containting messages such as "LETTER OF INTENT TO SUE" with a link to a compromised website. The website performs a drive-by-hacking comprising the viewing machine. Users are advised to be careful when following any links from emails. If you are unsure please forward the email to

September 30th, 2013

Broadband - DDoS traffic subsided at approximately 13:15. No further issues are anticipated.

Broadband - We have not yet received notification from the engineers, however packet loss appears to be reduced. We will post updates as and when they become available.

Broadband - We are currently seeing an issue with some connections that are experiencing packet loss/ intermittent connectivity. This appears to be a DDOS directed at parts of our DSL network. Engineers are working on resolving the issue and further updates will follow as soon as they become available.

September 15th, 2013

Broadband - Service is Online.

Virtual Servers - Service is Online.

Hosting - Service is Online.

Email - Service is Online.